# The Superconducting Quasiparticle-Amplifying Transmon: A Qubit-Based Sensor for meV Scale Phonons and Single THz Photons

### v3

@@ C.4 Signal Gain from Multiple Tunneling Events
$\tau_{qp\to JJ}\approx\frac{4\text{V}_{\text{trap}}}{\ell_{\text{JJ}}v_{QP}}$ (58)
@@ C.5 Total Quasiparticle Collection Efficiency
The final fraction of collected QPs starting with a stable QP population in the island is thus given by the product of Eqs. [50] [56] [57] [59]
@@
Figure 9: Simulated phonon energy resolution for single QP sensitivity as a function of island and trap length assuming negligible change in fill factor as a function of island length. Shown in the hatched region is twice the superconducting gap of Al, and the blacked out regions represent non-physical design parameters. Top, Resolution as a function of $p_{\text{trap}}$ with fixed $p_{\text{tunnel}}=1\times 10^{-6}$ for $p_{\text{trap}}=[1\times 10^{-3},1\times 10^{-4},1\times 10^{-5}]$ from left to right. Bottom, Resolution as a function of $p_{\text{tunnel}}$ with fixed $p_{\text{trap}}=1\times 10^{-4}$ for $p_{\text{tunnel}}=[5\times 10^{-7},1\times 10^{-6},5\times 10^{-6}]$ from left to right.
@@ Appendix D Sensor Bandwidth
While in previous sections we have focused on the low energy threshold of the SQUAT without considering the bandwidth of the readout method, the readout bandwidth will play the dominant role in limiting the the dynamic range of the sensor. The QP tunneling rate in a junction depends on the number of QPs near the junction, and since the number of QPs produced in the SQUAT depends on the initial energy of the particle, $E$, the measurement efficiency for higher energy events will decrease as the QP tunneling rate becomes greater than the readout rate, $1/2\tau_{\text{BW}}$. In this appendix we estimate the expected dynamic range of the SQUATs as a function of readout bandwidth.

### v4

@@ C.4 Signal Gain from Multiple Tunneling Events
$\tau_{qp\to JJ}\approx\frac{4\text{V}_{\text{trap}}}{\ell_{\text{JJ}}^{2}v_{QP}}$ (67)
@@ C.5 Total Quasiparticle Collection Efficiency
The final fraction of ‘collected QPs’ (where a collected QP is defined as a tunneling event) starting with a stable QP population in the island is thus given by the product of Eqs. [56] [61] [66] [68]
@@
Figure 11: Simulated phonon energy resolution for single QP sensitivity as a function of island and trap length assuming negligible change in fill factor as a function of island length. Shown in the hatched region is twice the superconducting gap of Al, and the blacked out regions represent non-physical design parameters. Top, Resolution as a function of $p_{\text{trap}}$ with fixed $p_{\text{tunnel}}=1\times 10^{-6}$ for $p_{\text{trap}}=[1\times 10^{-3},1\times 10^{-4},1\times 10^{-5}]$ from left to right. Bottom, Resolution as a function of $p_{\text{tunnel}}$ with fixed $p_{\text{trap}}=1\times 10^{-4}$ for $p_{\text{tunnel}}=[5\times 10^{-6},1\times 10^{-6},5\times 10^{-7}]$ from left to right.
@@ Appendix D Sensor Bandwidth
While in previous sections we have focused on only the low energy threshold of the SQUAT, we expect the dynamic range of the sensor to be limited primarily by the readout bandwidth. This can be understood intuitively as follows. The number of QPs produced in the SQUAT will scale linearly with the total energy $E$ of the original photon/phonon. When a large amount of QPs are near the junction, the average time between individual QP tunneling events will scale with the total number of QPs. Consequently, this means that the tunneling rate is a function of particle energy. In a practical readout mode, the SQUAT will have a finite rate a which it can make a measurement. This finite readout bandwidth implies that sequential tunneling events that take place faster than the readout rate of $1/2\tau_{\text{BW}}$ will likely not be measured. As such, the total number of QPs measured will no longer be linear with the original energy above a certain threshold, as set by the readout bandwidth. In this appendix we now derive where this threshold is, and how it sets the dynamic range of the SQUAT.